Given data is that car has a model that states use equation (7 at top 0 and bottom of integral) 7^0∫v(t)dt, 0-28m/s of 2.6s, time of reach 400m of 10.46s and tmaxspeed of 7s. I have calculated that value of coefficient A is 90.25m/s and time to reach max acceleration is 12.9m/s². I'm certain that the maximum velocity is V(t → ∞) = 90.25m/s. Please draw a graph for velocity v time showing velocity m/s on y axis also labelled with A, max velocity/speed. Then on the x axis please show where tmax is hit and where t400 is hit with x axis labelled time (s) please also show Vmax along the graph. once this is done please work out the areas for both sections and write a brief explanatation on the comparison between the model and my results :)
1
Expert's answer
2020-03-11T11:03:53-0400
"v(t)=90.25(1-e^{-t\/7})""v(2.6)=90.25(1-e^{-2.6/7})\approx28 (m/s)""v(7)=90.25(1-e^{-7/7})\approx57 (m/s)"




"s=\displaystyle\int_{0}^{9.675}v(t)dt=\displaystyle\int_{0}^{9.675}90.25(1-e^{-t/7})dt=""=90.25\bigg[t+7e^{-t/7}\bigg]\begin{matrix}
 9.675\\
 0
\end{matrix}=""=90.25(9.675+7e^{-9.675/7}-0-7)\approx400.01 (m)"

"v(9.675)=90.25(1-e^{-9.675/7})\approx67.59 (m/s)"


Check the work of the model amd compare the results


"v(2.6)=90.25(1-e^{-2.6/7})\approx28 (m/s), \ True"


"v(7)=90.25(1-e^{-7/7})\approx57 (m/s),\ True"

Time of reach 400m


"t_{400}=9.675\ s<10.46\ s"


"\delta={|9.675-10.46| \over 10.46}\cdot100\%\approx7.5\%"

I think based on the results of calculations, a model can be adopted to solve the problem.